OpenStudy (anonymous):

Game wardens use experiments to help determine the number of fish in a lake. Suppose 80 fish are caught,tagged, and released back into the lake. Two weeks later 120 fish are caught, 3 of which are found to have tags. Using this information, estimate the number of fish in the lake.

OpenStudy (anonymous):

3200 600 120 280

OpenStudy (sasogeek):

well if the first 80 caught were tagged and only 3 of that number were caught out of the 120 the second time... I don't know this for sure if it's a formula or anything, but logically I'd say that to estimate the total number of fishes in the lake, for every 120 fishes caught, there'll be 3 fishes that are tagged, distribute that over 80 tagged fishes and you have your estimated number of fishes. (80/3) * 120

OpenStudy (sasogeek):

@ganeshie8

OpenStudy (anonymous):

So would it be the first one?

OpenStudy (anonymous):

I'm not very good in math sorry

OpenStudy (sasogeek):

Like I said, I'm not sure myself, but year I'd go for the first one.

OpenStudy (sasogeek):

@myininaya

OpenStudy (anonymous):

@adajiamcneal @DARTHVADER2900 @iGreen @itsthatguy @KendrickLamar2014

OpenStudy (adajiamcneal):

the answer is 3,200 because 120 times 80 = 9,600 divided by 3 equals 3,200

mathslover (mathslover):

Let us keep it straight and simple. Use ratio : original caught / total = tagged caught (II time) / total caught (II time) 80 / x = 3 / 120 x = 80 * 120 / 3 x = 3200 So, A looks correct.

OpenStudy (anonymous):

Yay! Thanks everyone!

OpenStudy (adajiamcneal):

np